# Mitigating Omitted Variable Bias in Empirical Software Engineering

### v3

@@ 2.4 Other Forms of Confounding
The term “confounding” is used to denote different, related concepts in the statistical analysis of empirical data [26]. In this paper, we use “confounding” to denote bias in the estimate of a causal effect—as we demonstrated in a nutshell in the previous sections. This notion of confounding is customary in modern causal analysis [58], and in related approaches to mitigate confounding, such as instrumental variables [3].
An early usage of “confounding” in statistics was to denote noncollapsibility [26, 73]. In a nutshell, the effect of $X$ over $Y$ is collapsible over $Z$ if the marginal association between $X$ and $Y$ (obtained by averaging over $Z$) is the same as their conditional association (obtained by conditioning on $Z$); noncollapsibility holds when the two associations differ. While noncollapsibility is a purely statistical notion, the notion of confounding is rooted in causal relations, which depend not only on data but also on the data generation process [58]; therefore, the two concepts overlap but are distinct. In particular, noncollapsibility may signal causal confounding (for example, in a scenario like 1(c)); but it may also arise from nonlinearities without confounding. For example, consider a scenario with the same causal structure $X\to Y\leftarrow Z$ as in 1(b) but where $Y$ has a non-linear logistic relation with the linear combination of $X$ and $Z$. Since the causal structure is the same as in 1(b) there is no confounding; however, the effect of $X$ on $Y$ depends on whether we condition on $Z$, and hence there is noncollapsibility.
In classical frequentist statistics, the term “confounding” denotes an experimental design that makes two effects indistinguishable from data [20]. This usage of the term refers to statistical identifiability, and does not have, in general, any relation to causal effects. For example, a so-called fractional factorial design omits some combinations of factor levels, which may introduce confounding in this sense.
In fields such as psychometrics, confounding is often characterized as a measurement problem or, more generally, an issue of experimental design [15]. In the present paper, in contrast, we take the main point of view of analyzing observational data, having little or no control on the process that generated the data.
@@ 2.5 Time-dependent Data
As we explained in Section 1.2, the techniques presented in this paper are not directly applicable to analyze confounding in time-dependent data, such as those that are collected in longitudinal studies. This section is a brief detour into approaches that take time into account.
@@
Even though it is common for the observational data collected in software engineering studies (in particular by mining software repositories) to have a temporal dimension, a study may choose to abstract away the temporal dependencies and perform a cross-sectional analysis—possibly as a stepping stone towards a full-fledged longitudinal analysis. In the aforementioned domain of code reviews, a cross-sectional study would simply consider each data point as a “snapshot” of the relations $R_{k}\to Q_{k}$ at release (time) $k$. Imai and Kim [36] elucidate the impact of abstracting away time in longitudinal data, and under what assumptions it does not introduce confounding.
On the other hand, there have been several proposals in the literature of techniques to model and correct for confounding in time-dependent data. Marginal structural models (MSMs) [62] and structural nested models [76] are especially interesting from our perspective, since they are built atop the same causal DAGs that underlie this paper’s technique. For instance, MSMs build DAGs that represent the relation between variables at different time steps. In the code review domain, for example, we could posit a causal relation $E_{k}\to E_{k+1}$ if a reviewer may opt in to be assigned to the next code review.
@@ 3 Confounding in Programming Languages and Code Quality
Our first illustrative example follows closely the fundamental structure of Section 2.3’s prototypical example, while recasting it in a more realistic setting. Our goal is estimating the effect of using different programming languages (predictor variable Language) on the quality of the produced code written in that language (outcome variable Quality).
@@
As we discussed in depth in related work [21], [^7] In this section, we analyze a subset of the same data collected by others [60, 6] that we also analyzed in our previous work [21]. This section’s and our previous work’s analysis [21] are, however, orthogonal: in this section, we analyze possible confounding effects based on causal assumptions; in contrast, [21]’s analysis is purely statistical and targets Bayesian statistical modeling practices. a programmer’s ability (expressed by variable Skill) is likely to confound the causal effect of Language on Quality. [^8] It’s plausible several other confounders of this causal relation exist [21]; for clarity of presentation, we only consider Skills, as if it captures the effects of other possible confounders of the same relation. Namely, a more skilled programmer is likely to produce higher-quality code ($\textnormal{{Skill}}\to\textnormal{{Quality}}$); and programmers with different abilities may prefer to work with certain programming languages over others ($\textnormal{{Skill}}\to\textnormal{{Language}}$). Figure 3 summarizes these relations by means of a DAG, which is isomorphic [^9] Figure 3 and 1(c) are isomorphic as graphs. However, the variables associated with corresponding nodes may have different characteristics—most notably, $X$ is continuous whereas Language is categorical. to 1(c)’s abstract DAG.
As discussed in Section 2.3, this DAG structure entails that, if we want to estimate the true effect of Language on Quality for observational data—where we cannot control the effect of Skill on Language, that is, we cannot randomize which language each programmer will use—we need to also condition on the confounder Skill. In practice, this may be impossible because precisely measuring Skill is not easy: for example, if the data comes from a source code repository, the identity of the programmers may be unknown; even if we have access to a programmer’s identity, it may be practically cumbersome to reliably assess their programming skills.
@@ 3.1 Data: Programming Languages and Code Quality
To get a plausible ground truth about the relative effects of skills and languages on code quality, we analyze a subset of the data collected for a large-scale repository study [60], as made available in Berger et al. [6]’s reanalysis. This dataset summarizes the commit history of hundreds of projects in various languages.
To keep things simple—and to avoid bias that may come from underrepresented or misclassified languages [6]—we only retain data about projects: 1. written in Python or Java (two widely used, yet fairly different languages); 2. with at least 100 commits on record; 3. that are not multi-language (that is, each project is entirely written in Java, or entirely written in Python). These criteria select 105 projects: 45 written in Java and 60 written in Python. For each of these projects, we are interested in two key variables:
@@ 3.2 Quantifying Confounding
In this exercise, the goal is estimating the effect of choosing Java or Python as a programming language (variable Language) on the quality of the developed project (variable Quality). If we had access also to variable Skill, we could single out the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect by fitting 4(a)’s regression model $m_{3}$, which uses Language and Skill as predictors. On the data described in Section 3.1, this produces an estimate of the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ of $\ell=-0.012$, which indicates that using Python is very weakly associated with a modest reduction of quality (with a lot of uncertainty, as shown in 2(a)).
@@
In practice, it may not be possible to reliably measure the confounder Skill. In this case, we can only fit 4(b)’s model $m_{4}$, which gives us a different estimate $\ell=-0.052$ of the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ of Language on Quality. If we compare this with the previous estimate based on the unbiased model $m_{3}$, we notice that the confounding of Skill inflates the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ as measured in this data. In reality, we would not have access to the unbiased estimate; how to assess how much confidence we can put in an estimate that comes from a possibly confounded model? There are three main ways of proceeding [46]:
  • If we can muster an estimate the confounding effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$, we can use it to compute how strong the effect $\textnormal{{Skill}}\to\textnormal{{Language}}$ would have to be to tip the estimate of the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect (i.e., flip it from negative to positive). Section 3.2.2 discusses this scenario.
  • Conversely, if we can estimate the confounding effect $\textnormal{{Skill}}\to\textnormal{{Language}}$, we can compute how strong the effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ would have to be to tip the estimate of the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect. Section 3.2.3 discusses this scenario.
  • Finally, assuming that there are several different confounders that simultaneously affect Language and Quality, we can compute how many confounders with a certain effect would it take to tip the estimate of the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect. Section 3.2.4 discusses this scenario.
@@
To perform these so-called tipping point sensitivity analyses we will use the R package tipr [52], which implements state-of-the-art analysis techniques [46] based on causal models similar to those outlined in Section 2.2.
@@ 3.2.1 Scaled-mean Difference
In a tipping-point sensitivity analysis [52], any unmeasured confounder $Z$ is modeled as a standardized, normally distributed random variable. For a binary treatment variable $X$ with two nominal values $X_{0},X_{1}$, this means that the impact of the confounder $Z$ on the treatment $X$ is captured by normal distributions with unit variance and mean $\mu_{0}$ (when $X=X_{0}$) and $\mu_{1}$ (when $X=X_{1}$).
Accordingly, we use the scaled-mean difference (scd) to quantifying the impact of a confounder (Skill in our example) on the treatment (Language in our example) in a sensitivity analysis [52]. [^11] The scaled-mean difference $\mu_{1}/\sigma_{1}-\mu_{2}/\sigma_{2}$ between two groups is similar to, but different from, the standardized mean difference $(\mu_{1}-\mu_{2})/\sigma$—a commonly used effect size. Precisely, the scaled-mean difference is simply the difference of standardized means in the treatment and control groups. In our case, the $\textnormal{{scd}}(\textnormal{{Skill}}\to\textnormal{{Language}})$ of a programmer’s skills in different language is the difference between the mean Skill of a Python programmer and the mean Skill of a Java programmer, expressed as a multiple of Skill’s standard deviation:
@@ 3.2.4 Number of Confounders
Figure 5: A graphical summary of tipping analyses of confounder effect of Skill on the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect. For different measured values of effect ($\beta_{\textnormal{{Language}}\to\textnormal{{Quality}}}$), the plot shows the values of the effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ ($\beta_{\textnormal{{Skill}}\to\textnormal{{Quality}}}$) and of the scd $\textnormal{{Skill}}\to\textnormal{{Language}}$ ($\textsf{smd}_{\textnormal{{Skill}}\to\textnormal{{Language}}}$) that would tip the main effect. The two blue cross marks correspond to the two scenarios discussed in Section 3.2.2 and Section 3.2.3.
@@ 3.3 Sensitivity Analysis
Let’s generalize the tipping analysis described in this section beyond Ray et al. [60]’s data. First of all, consider a range of possible measured effects $\textnormal{{Language}}\to\textnormal{{Quality}}$, including both positive and negative values. For each of them, Figure 5 plots a line of a different color that marks the combination of values of confounding effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ (horizontal axis) and scd $\textnormal{{Skill}}\to\textnormal{{Language}}$ (vertical axis) that would tip the measured effect.
@@ 4 Confounding in Teamwork Effort
Our second illustrative example is based on Feldt et al. [18]’s review of research in the factors that affect the productivity of software development teams. It demonstrates how to practically assess the impact of potential unknown confounders in a more complex, realistic scenario. As in Section 3.2’s illustrative example, our aim is not to (directly) contribute to the knowledge about developer productivity; in fact, our analysis will at times adopt simplifying assumptions that are not (entirely) realistic. In contrast, the goal is to demonstrate credible confounding patterns, and to illustrate how to navigate around them in realistic conditions.
@@
Figure 6: A DAG summarizing relations among variables that characterize a software project, based on Feldt et al. [18]’s literature review. The variables are, from top to bottom and left to right: software Size, Team size, software Kind, Domain, Business area, Effort, Hardware, Location, Programming language, and Organization type.
@@ 4.1 Data: Teamwork, Effort, and Other Covariates
Feldt et al. [18] propose to use causal DAGs to summarize and combine the key findings of systematic literature reviews. In one of their case studies, they review several primary studies about productivity in software development. The DAG shown in Figure 6 is one of the DAGs that is obtained by applying Feldt et al. [18]’s approach; it summarizes the main relevant relations between variables that have been observed in some of the reviewed literature on the topic. [^12] This does not mean that this is the ultimate summary of research in the area of software development productivity. For our purposes, all that matters is that it displays a rich collection of plausible relations, so that our omitted variable bias analysis is grounded in a realistic scenario.
@@ 4.2 Adjustment Sets
Thus, addressing question i) above (What variables ought to be measured?) is tantamount to deciding which variables (covariates) [^13] In common statistical jargon, a covariate is any predictor variable other than the treatment. should be included in set $A$ in model $m_{A}$ (2). In particular, $A$ should include all confounders, so that the effect $\textnormal{{T}}\to\textnormal{{E}}$ can be estimated without bias. Set $A$ represents what is called an adjustment set: given a DAG and two of its variables—representing the outcome (E in our case) and the treatment (T in our case)—the adjustment set $A$ is the set of additional variables in the DAG that we have to include in (2)’s model to ensure that $\beta$ estimates the unbiased, uncounfounded, genuine causal effect of T on E. In other words, the adjustment set includes all predictors that we should include if we want to avoid introducing omitted variable bias.
The adjustment set (or rather adjustment sets, since a DAG may admit multiple, alternative adjustment sets) can be computed directly on the DAG based solely on its structure—assuming, of course, that the DAG correctly captures real-world causal relations. [^14] In previous work [22], we illustrated how adjustment sets are computed, and what’s the intuition behind them. For brevity, we do not repeat the explanation in this paper, but simply use the adjustment sets computed from a DAG in our analysis. In our example, the adjustment set is $A=\{\textnormal{{O}},\textnormal{{S}}\}$; therefore including the two predictors O and S, in addition to T, ensures that $\beta$ captures the net causal relation between treatment and outcome.
Remember that the adjustment set is a variable selection strategy whose goal is correcting for possible confounding effects of causal relations. As explained in Section 1.2, this is the specific focus of this paper; naturally, different kinds of research questions may require different variable selection criteria. [^15] In particular, if the goal is fitting a statistical (regression) model with an optimal trade off between predictive accuracy and model conciseness, information-theoretic model selection criteria provide a serviceable approach [16, 80, 77]. This does not necessarily imply that adjusting for confounding bias is irreconcilable with other data analysis goals. In fact, we will see that a DAG often admits different equivalent adjustment sets, which often provides more flexibility in how this technique is applied.

### v4

@@ 2.4 Other Forms of Confounding
The term “confounding” is used to denote different, related concepts in the statistical analysis of empirical data (Greenland and Morgenstern, 2001). In this paper, we use “confounding” to denote bias in the estimate of a causal effect—as we demonstrated in a nutshell in the previous sections. This notion of confounding is customary in modern causal analysis (Pearl, 2009), and in related approaches to mitigate confounding, such as instrumental variables (Angrist et al., 1996).
An early usage of “confounding” in statistics was to denote noncollapsibility (Greenland and Morgenstern, 2001; Undy Yule, 1903). In a nutshell, the effect of $X$ over $Y$ is collapsible over $Z$ if the marginal association between $X$ and $Y$ (obtained by averaging over $Z$) is the same as their conditional association (obtained by conditioning on $Z$); noncollapsibility holds when the two associations differ. While noncollapsibility is a purely statistical notion, the notion of confounding is rooted in causal relations, which depend not only on data but also on the data generation process (Pearl, 2009); therefore, the two concepts overlap but are distinct. In particular, noncollapsibility may signal causal confounding (for example, in a scenario like 1(c)); but it may also arise from nonlinearities without confounding. For example, consider a scenario with the same causal structure $X\to Y\leftarrow Z$ as in 1(b) but where $Y$ has a non-linear logistic relation with the linear combination of $X$ and $Z$. Since the causal structure is the same as in 1(b) there is no confounding; however, the effect of $X$ on $Y$ depends on whether we condition on $Z$, and hence there is noncollapsibility.
In classical frequentist statistics, the term “confounding” denotes an experimental design that makes two effects indistinguishable from data (Fisher, 1935). This usage of the term refers to statistical identifiability, and does not have, in general, any relation to causal effects. For example, a so-called fractional factorial design omits some combinations of factor levels, which may introduce confounding in this sense.
In fields such as psychometrics, confounding is often characterized as a measurement problem or, more generally, an issue of experimental design (Cook and Campbell, 1979). In the present paper, in contrast, we take the main point of view of analyzing observational data, having little or no control on the process that generated the data.
@@ 2.5 Time-dependent Data
As we explained in subsection 1.2, the techniques presented in this paper are not directly applicable to analyze confounding in time-dependent data, such as those that are collected in longitudinal studies. This section is a brief detour into approaches that take time into account.
@@
Even though it is common for the observational data collected in software engineering studies (in particular by mining software repositories) to have a temporal dimension, a study may choose to abstract away the temporal dependencies and perform a cross-sectional analysis—possibly as a stepping stone towards a full-fledged longitudinal analysis. In the aforementioned domain of code reviews, a cross-sectional study would simply consider each data point as a “snapshot” of the relations $R_{k}\to Q_{k}$ at release (time) $k$. Imai and Kim (2019) elucidate the impact of abstracting away time in longitudinal data, and under what assumptions it does not introduce confounding.
On the other hand, there have been several proposals in the literature of techniques to model and correct for confounding in time-dependent data. Marginal structural models (MSMs) Robins et al. (2000) and structural nested models Vansteelandt and Joffe (2014) are especially interesting from our perspective, since they are built atop the same causal DAGs that underlie this paper’s technique. For instance, MSMs build DAGs that represent the relation between variables at different time steps. In the code review domain, for example, we could posit a causal relation $E_{k}\to E_{k+1}$ if a reviewer may opt in to be assigned to the next code review.
@@ 3 Confounding in Programming Languages and Code Quality
Our first illustrative example follows closely the fundamental structure of subsection 2.3’s prototypical example, while recasting it in a more realistic setting. Our goal is estimating the effect of using different programming languages (predictor variable Language) on the quality of the produced code written in that language (outcome variable Quality).
@@
As we discussed in depth in related work (Furia et al., 2022), [^7] In this section, we analyze a subset of the same data collected by others (Ray et al., 2014; Berger et al., 2019) that we also analyzed in our previous work (Furia et al., 2022). This section’s and our previous work’s analysis (Furia et al., 2022) are, however, orthogonal: in this section, we analyze possible confounding effects based on causal assumptions; in contrast, (Furia et al., 2022)’s analysis is purely statistical and targets Bayesian statistical modeling practices. a programmer’s ability (expressed by variable Skill) is likely to confound the causal effect of Language on Quality. [^8] It’s plausible several other confounders of this causal relation exist (Furia et al., 2022); for clarity of presentation, we only consider Skills, as if it captures the effects of other possible confounders of the same relation. Namely, a more skilled programmer is likely to produce higher-quality code ($\textnormal{{Skill}}\to\textnormal{{Quality}}$); and programmers with different abilities may prefer to work with certain programming languages over others ($\textnormal{{Skill}}\to\textnormal{{Language}}$). Figure 3 summarizes these relations by means of a DAG, which is isomorphic [^9] Figure 3 and 1(c) are isomorphic as graphs. However, the variables associated with corresponding nodes may have different characteristics—most notably, $X$ is continuous whereas Language is categorical. to 1(c)’s abstract DAG.
As discussed in subsection 2.3, this DAG structure entails that, if we want to estimate the true effect of Language on Quality for observational data—where we cannot control the effect of Skill on Language, that is, we cannot randomize which language each programmer will use—we need to also condition on the confounder Skill. In practice, this may be impossible because precisely measuring Skill is not easy: for example, if the data comes from a source code repository, the identity of the programmers may be unknown; even if we have access to a programmer’s identity, it may be practically cumbersome to reliably assess their programming skills.
@@ 3.1 Data: Programming Languages and Code Quality
To get a plausible ground truth about the relative effects of skills and languages on code quality, we analyze a subset of the data collected for a large-scale repository study (Ray et al., 2014), as made available in Berger et al. (2019)’s reanalysis. This dataset summarizes the commit history of hundreds of projects in various languages.
To keep things simple—and to avoid bias that may come from underrepresented or misclassified languages (Berger et al., 2019)—we only retain data about projects: 1. written in Python or Java (two widely used, yet fairly different languages); 2. with at least 100 commits on record; 3. that are not multi-language (that is, each project is entirely written in Java, or entirely written in Python). These criteria select 105 projects: 45 written in Java and 60 written in Python. For each of these projects, we are interested in two key variables:
@@ 3.2 Quantifying Confounding
In this exercise, the goal is estimating the effect of choosing Java or Python as a programming language (variable Language) on the quality of the developed project (variable Quality). If we had access also to variable Skill, we could single out the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect by fitting 4(a)’s regression model $m_{3}$, which uses Language and Skill as predictors. On the data described in subsection 3.1, this produces an estimate of the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ of $\ell=-0.012$, which indicates that using Python is very weakly associated with a modest reduction of quality (with a lot of uncertainty, as shown in 2(a)).
@@
In practice, it may not be possible to reliably measure the confounder Skill. In this case, we can only fit 4(b)’s model $m_{4}$, which gives us a different estimate $\ell=-0.052$ of the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ of Language on Quality. If we compare this with the previous estimate based on the unbiased model $m_{3}$, we notice that the confounding of Skill inflates the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ as measured in this data. In reality, we would not have access to the unbiased estimate; how to assess how much confidence we can put in an estimate that comes from a possibly confounded model? There are three main ways of proceeding (Lin et al., 1998):
  • If we can muster an estimate the confounding effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$, we can use it to compute how strong the effect $\textnormal{{Skill}}\to\textnormal{{Language}}$ would have to be to tip the estimate of the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect (i.e., flip it from negative to positive). subsubsection 3.2.2 discusses this scenario.
  • Conversely, if we can estimate the confounding effect $\textnormal{{Skill}}\to\textnormal{{Language}}$, we can compute how strong the effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ would have to be to tip the estimate of the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect. subsubsection 3.2.3 discusses this scenario.
  • Finally, assuming that there are several different confounders that simultaneously affect Language and Quality, we can compute how many confounders with a certain effect would it take to tip the estimate of the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect. subsubsection 3.2.4 discusses this scenario.
@@
To perform these so-called tipping point sensitivity analyses we will use the R package tipr (McGowan, 2022), which implements state-of-the-art analysis techniques (Lin et al., 1998) based on causal models similar to those outlined in subsection 2.2.
@@ 3.2.1 Scaled-mean Difference
In a tipping-point sensitivity analysis (McGowan, 2022), any unmeasured confounder $Z$ is modeled as a standardized, normally distributed random variable. For a binary treatment variable $X$ with two nominal values $X_{0},X_{1}$, this means that the impact of the confounder $Z$ on the treatment $X$ is captured by normal distributions with unit variance and mean $\mu_{0}$ (when $X=X_{0}$) and $\mu_{1}$ (when $X=X_{1}$).
Accordingly, we use the scaled-mean difference (scd) to quantifying the impact of a confounder (Skill in our example) on the treatment (Language in our example) in a sensitivity analysis (McGowan, 2022). [^11] The scaled-mean difference $\mu_{1}/\sigma_{1}-\mu_{2}/\sigma_{2}$ between two groups is similar to, but different from, the standardized mean difference $(\mu_{1}-\mu_{2})/\sigma$—a commonly used effect size. Precisely, the scaled-mean difference is simply the difference of standardized means in the treatment and control groups. In our case, the $\textnormal{{scd}}(\textnormal{{Skill}}\to\textnormal{{Language}})$ of a programmer’s skills in different language is the difference between the mean Skill of a Python programmer and the mean Skill of a Java programmer, expressed as a multiple of Skill’s standard deviation:
@@ 3.2.4 Number of Confounders
Figure 5: A graphical summary of tipping analyses of confounder effect of Skill on the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect. For different measured values of effect ($\beta_{\textnormal{{Language}}\to\textnormal{{Quality}}}$), the plot shows the values of the effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ ($\beta_{\textnormal{{Skill}}\to\textnormal{{Quality}}}$) and of the scd $\textnormal{{Skill}}\to\textnormal{{Language}}$ ($\textsf{smd}_{\textnormal{{Skill}}\to\textnormal{{Language}}}$) that would tip the main effect. The two blue cross marks correspond to the two scenarios discussed in subsubsection 3.2.2 and subsubsection 3.2.3.
@@ 3.3 Sensitivity Analysis
Let’s generalize the tipping analysis described in this section beyond Ray et al. (2014)’s data. First of all, consider a range of possible measured effects $\textnormal{{Language}}\to\textnormal{{Quality}}$, including both positive and negative values. For each of them, Figure 5 plots a line of a different color that marks the combination of values of confounding effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ (horizontal axis) and scd $\textnormal{{Skill}}\to\textnormal{{Language}}$ (vertical axis) that would tip the measured effect.
@@ 4 Confounding in Teamwork Effort
Our second illustrative example is based on Feldt et al. (2025)’s review of research in the factors that affect the productivity of software development teams. It demonstrates how to practically assess the impact of potential unknown confounders in a more complex, realistic scenario. As in subsection 3.2’s illustrative example, our aim is not to (directly) contribute to the knowledge about developer productivity; in fact, our analysis will at times adopt simplifying assumptions that are not (entirely) realistic. In contrast, the goal is to demonstrate credible confounding patterns, and to illustrate how to navigate around them in realistic conditions.
@@
Figure 6: A DAG summarizing relations among variables that characterize a software project, based on Feldt et al. (2025)’s literature review. The variables are, from top to bottom and left to right: software Size, Team size, software Kind, Domain, Business area, Effort, Hardware, Location, Programming language, and Organization type.
@@ 4.1 Data: Teamwork, Effort, and Other Covariates
Feldt et al. (2025) propose to use causal DAGs to summarize and combine the key findings of systematic literature reviews. In one of their case studies, they review several primary studies about productivity in software development. The DAG shown in Figure 6 is one of the DAGs that is obtained by applying Feldt et al. (2025)’s approach; it summarizes the main relevant relations between variables that have been observed in some of the reviewed literature on the topic. [^12] This does not mean that this is the ultimate summary of research in the area of software development productivity. For our purposes, all that matters is that it displays a rich collection of plausible relations, so that our omitted variable bias analysis is grounded in a realistic scenario.
@@ 4.2 Adjustment Sets
Thus, addressing question item i) above (What variables ought to be measured?) is tantamount to deciding which variables (covariates) [^13] In common statistical jargon, a covariate is any predictor variable other than the treatment. should be included in set $A$ in model $m_{A}$ (2). In particular, $A$ should include all confounders, so that the effect $\textnormal{{T}}\to\textnormal{{E}}$ can be estimated without bias. Set $A$ represents what is called an adjustment set: given a DAG and two of its variables—representing the outcome (E in our case) and the treatment (T in our case)—the adjustment set $A$ is the set of additional variables in the DAG that we have to include in (2)’s model to ensure that $\beta$ estimates the unbiased, uncounfounded, genuine causal effect of T on E. In other words, the adjustment set includes all predictors that we should include if we want to avoid introducing omitted variable bias.
The adjustment set (or rather adjustment sets, since a DAG may admit multiple, alternative adjustment sets) can be computed directly on the DAG based solely on its structure—assuming, of course, that the DAG correctly captures real-world causal relations. [^14] In previous work (Furia et al., 2023), we illustrated how adjustment sets are computed, and what’s the intuition behind them. For brevity, we do not repeat the explanation in this paper, but simply use the adjustment sets computed from a DAG in our analysis. In our example, the adjustment set is $A=\{\textnormal{{O}},\textnormal{{S}}\}$; therefore including the two predictors O and S, in addition to T, ensures that $\beta$ captures the net causal relation between treatment and outcome.
Remember that the adjustment set is a variable selection strategy whose goal is correcting for possible confounding effects of causal relations. As explained in subsection 1.2, this is the specific focus of this paper; naturally, different kinds of research questions may require different variable selection criteria. [^15] In particular, if the goal is fitting a statistical (regression) model with an optimal trade off between predictive accuracy and model conciseness, information-theoretic model selection criteria provide a serviceable approach (Dvorzak and Wagner, 2016; Watanabe, 2010; Vehtari et al., 2017). This does not necessarily imply that adjusting for confounding bias is irreconcilable with other data analysis goals. In fact, we will see that a DAG often admits different equivalent adjustment sets, which often provides more flexibility in how this technique is applied.